OpenStudy (anonymous):

how would you factor (a+b)^2 - 64

OpenStudy (xapproachesinfinity):

64 is perfect square

OpenStudy (xapproachesinfinity):

\(x^2-y^2=(x+y)(x-y)\) you will need this difference of two perfect squares

OpenStudy (xapproachesinfinity):

\((a+b)^2\) this quantity is a perfect square by it self

OpenStudy (xapproachesinfinity):

any clue? what is the number if we multiply it by it self we get 64

OpenStudy (anonymous):

yeah the number would be 8

OpenStudy (xapproachesinfinity):

Good so you agree that \(64=8*8=8^2\)

OpenStudy (anonymous):

yes i do agree so I would break it apart by (a+8)(a-8) because its negative 64

OpenStudy (xapproachesinfinity):

no no! just follow since you agreed now we can write \((a+b)^2-64=(a+b)^2-8^2\)

OpenStudy (xapproachesinfinity):

yes?

OpenStudy (anonymous):

ok yes i see that

OpenStudy (xapproachesinfinity):

now treat a+b as one entity let's call it z a+b=z alright so \((a+b)^2-8^2=z^2-8^2\)

OpenStudy (xapproachesinfinity):

i don't need to always call it some name i just want to explain it to you so i becomes easy to deal with okay

OpenStudy (anonymous):

yep no problem

OpenStudy (xapproachesinfinity):

now we have \(z^2-8^2\) difference of two square like is said before \(x^2-y^2=(x-y)(x+y)\) we do the same with \(z^2-8^2=(z-8)(z+8)\)

OpenStudy (xapproachesinfinity):

yes?!

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

now z=a+b because i just called z it was a+b so i have \(z^2-8^2=(z-8)(z+8)=(a+b-8)(a+b+8)\)

OpenStudy (xapproachesinfinity):

So \((a+b)^2-64=(a+b-8)(a+b+8)\)

OpenStudy (anonymous):

ok that makes so much more sense thank you for your help

OpenStudy (xapproachesinfinity):

My pleasure! once you know how to do it you don't need all those steps you go directly to write you factored form because we know that's the difference of two perfect squares

OpenStudy (anonymous):

yeah i understood the difference of perfect squares i was just confusing myself when i had to difference of perfect squares to combine

OpenStudy (xapproachesinfinity):

the only step you need is \((a+b)^2-64=(a+b)^2-8^2\) then you write what it is directly

OpenStudy (xapproachesinfinity):

if you foil this \((x+y)(x-y)\) you will get \(x^2-y^2\) which is called difference of perfect squares

OpenStudy (anonymous):

yeah i understand that now thank you do you think you can help with other problem too?

OpenStudy (xapproachesinfinity):

welcome! actually i have to go out now sorry

OpenStudy (anonymous):

ok no problem

OpenStudy (xapproachesinfinity):

oh if my help satisfied you there is best response thing click on it^_^